calculate the potential of hydrogen electrode whose ph=10

Grade:11

1 Answers

Akshay Meena
10 years ago
According to Nerst equation E = E0-2.303(RT/nF)logQ Here : E0 = 0 [H+]=10^-10, Q = 10^10 as Pressure of H2= 1 atm and n = 1 gives where R = 8.314, T = 298 K generally), F = 96500 C E = 0-2.303(RT/F)(10)=-23.03RT/F = -0.591 V, Thanks & Regards Akshay Meena, askIITians Faculty M.Tech, IIT Kharagpur